Circular No. 7825

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002ao, 2002ap, 2002ar, 2002au, 2002av A. V. Filippenko and R. Chornock, University of California, Berke-

ley, report that inspection of CCD spectra (range 330–1000 nm), obtained on Feb. 11 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2002ap (IAUC 7810) is very similar to the peculiar type-Ic SN 1997ef (‘hypernova’), ∼ 2 weeks after the explosion (Iwamoto et al. 2000, Ap.J. 534, 660, Fig. 1; Matheson et al. 2001, A.J. 121, 1648, Fig. 18), as was suggested by Kinugasa et al. (IAUC 7811). Broad emission emission features are seen near 460 nm, 535 nm, and 645 nm. Absorption troughs nominally associated with O I 777.4-nm and the Ca II triplet are visible near 755 nm and 810 nm, respectively; the expansion velocity derived from the O I line is ∼ 9000 km/s. SN 2002av (IAUC 7823) is of type Ia near maximum brightness. SN 2002ar (IAUC 7819) is also of type Ia, with a spectral-feature age of 1 ± 2 days before maximum. SN 2002au (IAUC 7823) is a supernova, probably (but not definitely) of type Ia, ∼ 1–2 weeks past maximum. SN 2002ao (IAUC 7809), whose published classification is uncertain (IAUC 7810, 7815, 7819), has a very unusual spectrum dominated by strong He I emission lines with FWHM ∼ 3000–4000 km/s. The corresponding hydrogen emission is very weak, but broader lines of O I 777.4-nm (FWHM = 6000 km/s), the Ca II triplet, and Fe II are present. There is a striking resemblance to the spectrum of SN 1999cq (Matheson et al. 2000, A.J. 119, 2303).

A. Riffeser, C. A. Goessl, and C. Ries, Institute of Astronomy and Astrophysics, Munich, report CCD magnitudes for SN 2002ap: Feb. 1.8 UT, U = 13.48, B = 13.46, R = 12.88; 2.8, 13.35, 13.30, 12.79; 3.8, 13.29, 13.20, 12.65; 4.8, 13.26, 13.13, 12.54; 5.8, 13.34, 13.10, 12.42. A 300-s exposure taken on 1999 Sept. 9 shows no progenitor to limiting mag R = 21.1 (3σ), confirming the report by Vreeswijk et al. on IAUC 7820.

NOVAE IN M31 Filippenko and Chornock also write: “Inspection of CCD spectra, ob-

tained on Feb. 11 UT as above, reveals that the first two nova candidates listed on IAUC 7794 are indeed novae, with strong Hα emission. (The third candidate was not observed.) The FWHM of Hα is ∼ 3000 km/s in the first nova and 1500 km/s in the second nova, although in both cases the profile is rather boxlike.” It should be noted that, unlike the KAIT offsets, the Padua east-west offsets given on IAUC 7794 did not include the cos δ factor.

2002 February 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green